find the shortest distance between the origin and the surface x ^ 2y-z ^ 2 + 9 = 0
1
Expert's answer
2020-04-30T19:25:40-0400

Let P(x,y,z) be the point on the surface "x^{2}y-z^2+9=0" . Then the distance from the origin is "d = \sqrt{x^2+y^2+z^2}" . We have to minimize the value of "d" subject to the constraint "x^{2}y-z^2+9=0".


Let "f(x,y,z) = x^{2}+y^{2}+z^{2}" and let "g(x,y,z)=x^{2}y-z^{2}+9".

From the given constraint, "z^{2}=x^{2}y+9". Using this in f(x,y,z) it becomes "x^{2}+y^{2}+x^{2}y+9", which is a function in two variables.


Let "F(x,y)=x^{2}+y^{2}+x^{2}y+9." Now we have to solve,

"\dfrac{\partial F}{\partial x} = 0;
\dfrac{\partial F}{\partial y} = 0\\"

which gives,

"2x+2xy = 0~~~~(1)\\
2y+x^{2}=0~~~~~~~(2)"


Solving equation (1), we get either "x=0~~\text{or}~~y=-1."


When "x = 0, y = 0" and When "y = -1, x = \pm \sqrt{2}" , from equation (2).

Now from the constraint equation, "z^{2} = 9 \Rightarrow z = \pm 3", when x = 0 and y=0.

Also, "z^{2} = -2 + 9 = 7 \Rightarrow z = \pm \sqrt{7}", when "y = -1, x = \pm \sqrt{2}".


Therefore the critical points are, "(0,0,\pm3), (\pm \sqrt{2},-1,\pm\sqrt{7})".

Now,

"f(0,0,\pm3) = 9\\
f(\pm\sqrt2,-1,\pm\sqrt7)=10"


Hence the minimum distance is "d = 3" which occurs at the points "(0,0,\pm3)".